OpenStudy (anonymous):

What does 2a+2b=c equal solving for b help please

OpenStudy (john_es):

Solving for b, \[2a+2b-2a=c-2a\Rightarrow2b=c-2a\Rightarrow b=\frac{c-2a}{2}\]